OpenStudy (oswaldmurphy):

I need help.. @phi or anyone else who is good wilth algebra 1.. Its an honors assignment.

OpenStudy (oswaldmurphy):

@pooja195 , look at this http://prntscr.com/bvalx7 Task 1.

OpenStudy (oswaldmurphy):

@Jadeishere @Hero @Data_Lg2 @sammixboo @pooja195

OpenStudy (data_lg2):

@OswaldMurphy are you still there? Let's start with \(\sf Task 1\) First, assign any values for a, b, c, and d. Can you do that? a= b= c= d=

OpenStudy (oswaldmurphy):

Yes, I think I got it.

OpenStudy (data_lg2):

would you mind telling me what you have done so far? (:

OpenStudy (oswaldmurphy):

Part 1 is f(x) = (x – 2)/4 g(x) = 4x + 2

OpenStudy (oswaldmurphy):

Part 2 is, y = (x - 2)/4. We find the inverse so, x = (y - 2)/4. Multiply by 4, 4x = y - 2. Add 2, y = 4x + 2. Ill brb, ok?

OpenStudy (oswaldmurphy):

y = cx - 2 = y = cx - 3 because of inverse.

OpenStudy (oswaldmurphy):

@Data_Lg2 @Jadeishere @ganeshie8

OpenStudy (oswaldmurphy):

I got this, Part 1). f(x) = (x – 2)/4 g(x) = 4x + 2 Part 2). y = (x - 2)/4. We find the inverse so, x = (y - 2)/4. Multiply by 4, 4x = y - 2. Add 2, y = 4x + 2. y = cx - 2 = y = cx - 3 because of inverse.

OpenStudy (data_lg2):

Sorry I had to refresh. So far you are going great. y = (x - 2)/4. Switching x & y variables: x = (y - 2)/4. Multiply by 4: 4x = y - 2 Add 2 on both sides: y = 4x + 2. ---> this function equals the function g(x) For part 3: You just replace the 'x' variable of the function g(x) with the function f(x) and then evaluate. \(\sf \Large g(f(x)) = 4( \color{red}{\frac{x-2}{4}})+2\)

OpenStudy (data_lg2):

tell me what you will get..

OpenStudy (oswaldmurphy):

How do i solve that 0-0

OpenStudy (data_lg2):

you just simplify it. Do this part first, what will you end up with? \(\sf \Large 4( \color{red}{\frac{x-2}{4}})\)

OpenStudy (oswaldmurphy):

x-2..

OpenStudy (data_lg2):

right, now you have \(\sf \Large g(f(x)) = \color{red}{x-2}+2\) simplify it more and you will get...?

OpenStudy (oswaldmurphy):

Sorry I got suspended -_-

OpenStudy (oswaldmurphy):

:O g(f(x)) = x OMG

OpenStudy (data_lg2):

yep, you got it.

OpenStudy (data_lg2):

I gtg, I think you can manage to do Part 4 now (: For task 2, try opening a new thread so that others can help you. Sorry, I have to go somewhere rn..